# Central Path Proximal Policy Optimization

### v1

@@ 4 Computational Experiments
To evaluate our approach, we conduct experiments aimed at testing the benefits of using central path approximation as a design principle for constrained policy optimization algorithms. We benchmark C3PO against a range of representative constrained reinforcement learning baselines. We include methods from three major algorithmic families: penalty-based methods (P3O, P2BPO), Lagrangian methods (PPO-Lag, CPPO-PID), and trust-region methods (CPO, C-TRPO).
Conceptually, penalty-based methods, especially algorithms that augment the PPO loss with a penalty, like P3O Zhang et al. (2022) and P2BPO Dey et al. (2024), are closest to our approach. Like C3PO, those penalize constraint violations directly in the policy gradient loss using a ReLU-penalty. Lagrangian methods maintain dual variables to enforce constraints adaptively. PPO-Lagrangian Ray et al. (2019) applies this principle to the PPO algorithm, forming a loss which is similar to C3PO’s. For completeness, we consider CPPO-PID Stooke et al. (2020) as a more recent Lagrangian baseline. Finally, trust region methods, such as CPO Achiam et al. (2017) and C-TRPO Milosevic et al. (2025), use trust regions and constrained updates to maintain stable reward improvement and feasibility throughout training. They do not aim for scalability, but form strong baselines on the considered benchmark.
We benchmark the algorithms on 4 locomotion tasks and 4 navigation tasks from Safety Gymnasium Ji et al. (2023), as in Milosevic et al. (2025). For the baseline algorithms, we use the hyper-parameters reported in Ji et al. (2023), and for P3O and C-TRPO the recommended parameters in Zhang et al. (2022) and Milosevic et al. (2025) respectively. For C3PO we use $\kappa=30.0$ and $w=0.05$. Each algorithm is trained on each task for 10 million steps with a cost threshold of 25.0. Final iterate performance is measured by aggregating over 5 seeds using rliable Agarwal et al. (2021).
@@
The results provide confirmatory evidence for the usefulness of the central path approach: policies trained with C3PO exhibit a stable progression toward the constrained optimum, maintaining feasibility for most training iterations, see Figure 4. Furthermore, C3PO consistently outperforms prior PPO-style penalty methods in terms of achieved reward, while also adhering more strictly to the specified constraints, see Figure 3. This improved trade-off between reward and feasibility offers additional support for the effectiveness of the central path approach. While C3PO does not outperform trust-region methods across all tasks in the benchmark, it performs well consistently, resulting in high aggregated performance. The full benchmark results table and more examples like Figure 4 are presented in Appendix C.
@@ 5 Conclusion
In this work, we use central path approximation as a guiding principle for designing policy optimization methods for constrained RL. We propose C3PO, an algorithm which is obtained through a simple augmentation of the original PPO-loss inspired by the central path approach. Our experimental results support this design principle: Compared to existing PPO-style penalty and Lagrangian methods, C3PO exhibits improved performance with tighter constraint satisfaction, highlighting the benefits of a central path approach in constrained policy optimization. While the current results are limited to small-scale simulations and simplified settings, such as a single constraint per task, they suggest that the central path-approximation is a promising design principle for constrained RL algorithms. We hope this early-stage contribution encourages further discussion and refinement of on-policy safe policy optimization algorithms and other methods that solve constrained MDPs. Future work will explore extensions to high-dimensional tasks, theoretical guarantees, and applications such as safety-critical control and LLM fine-tuning. We hope this early-stage contribution encourages further dialogue on improving constrained RL algorithms using insights from constrained optimization.
@@ B.1 Exact Penalty Methods
We provide a general result for the exactness of the penalties considered here for general discussions of exact penalty methods, we refer to the standard textbooks in optimization Bertsekas [1997], Nocedal and Wright [1999]. Here, we consider a compact subset $X\subset\mathbb{R}^{n}$ with non-empty interior a differentiable functions $f,g\in C^{1}(X)$ and the constrained optimization problem
@@ B.2 C3PO Exact Penalty
and for $\mathbb{A}_{\textup{B}}\geq\mathbb{A}_{c}^{\pi_{k}}(\pi)>0$ it holds that $\delta_{\textup{B}}\geq D_{\textup{B}}(\pi,\pi_{k})>0$. To solve for $\mathbb{A}_{\textup{B}}$, we rewrite 33 as
@@
$P_{\textup{Barrier}}$ =$=\{\pi:\mathbb{A}_{c}^{\pi_{k}}(\pi)-\mathbb{A}_{\textup{B}}<0,\ \mathbb{A}_{c}^{\pi_{k}}(\pi)\geq 0\},$ (37)
=$=\{\pi:\mathbb{A}_{c}^{\pi_{k}}(\pi)-b\ (W(-\exp(-\delta_{\textup{B}}-1))+1)<0,\ \mathbb{A}_{c}^{\pi_{k}}(\pi)\geq 0\},$ (38)
=$=\{\pi:\mathbb{A}_{c}^{\pi_{k}}(\pi)-b\ w<0,\ \mathbb{A}_{c}^{\pi_{k}}(\pi)\geq 0\},$ (39)
@@
[…]
@@
[…]

### v2

@@ 4 Computational Experiments
To evaluate our approach, we conduct experiments aimed at testing the benefits of using central path approximation as a design principle for constrained policy optimization algorithms. [^1] Code: https://github.com/milosen/c3po We benchmark C3PO against a range of representative constrained reinforcement learning baselines. We include methods from three major algorithmic families: penalty-based methods (P3O, P2BPO), Lagrangian methods (PPO-Lag, CPPO-PID), and trust-region methods (CPO, C-TRPO).
Conceptually, penalty-based methods, especially algorithms that augment the PPO loss with a penalty, like P3O (Zhang et al., 2022) and P2BPO (Dey et al., 2024), are closest to our approach. Like C3PO, those penalize constraint violations directly in the policy gradient loss using a ReLU-penalty. Lagrangian methods maintain dual variables to enforce constraints adaptively. PPO-Lagrangian (Ray et al., 2019) applies this principle to the PPO algorithm, forming a loss which is similar to C3PO’s. For completeness, we consider CPPO-PID (Stooke et al., 2020) as a more recent Lagrangian baseline. Finally, trust region methods, such as CPO (Achiam et al., 2017) and C-TRPO (Milosevic et al., 2025), use trust regions and constrained updates to maintain stable reward improvement and feasibility throughout training. They do not aim for scalability, but form strong baselines on the benchmarks.
We benchmark the algorithms on 4 locomotion tasks and 4 navigation tasks from Safety Gymnasium (Ji et al., 2023), as done by Milosevic et al. (2025). For the baseline algorithms, we use the hyper-parameters reported by Ji et al. (2023), and for P3O and C-TRPO the recommended parameters in Zhang et al. (2022) and Milosevic et al. (2025) respectively. For C3PO we use $\kappa=30.0$ and $w=0.05$. Each algorithm is trained on each task for 10 million steps with a cost threshold of 25.0. Final iterate performance is measured by aggregating over 5 seeds using rliable (Agarwal et al., 2021).
@@
The results provide confirmatory evidence for the usefulness of the central path approach. Policies trained with C3PO exhibit a stable progression toward the constrained optimum, maintaining feasibility for most training iterations, see Figure 4. Furthermore, C3PO consistently outperforms prior PPO-style penalty methods in terms of achieved reward, while also adhering more strictly to the specified constraints, see Figure 3. This improved trade-off between reward and feasibility offers additional support for the effectiveness of the central path approach. While C3PO does not outperform trust-region methods across all tasks in the benchmark, it performs well consistently, resulting in high aggregated performance. The full benchmark results table and more examples like Figure 4 are presented in Appendix C.
@@ 5 Conclusion
In this work, we use central path approximation as a guiding principle for designing policy optimization methods for constrained RL. We propose C3PO, an algorithm which is obtained through a simple augmentation of the original PPO-loss inspired by the central path approach. Our experimental results support this design principle: Compared to existing PPO-style penalty and Lagrangian methods, C3PO exhibits improved performance with tighter constraint satisfaction, highlighting the benefits of a central path approach in constrained policy optimization.
While the current results are limited to small-scale simulations and simplified settings, such as a single constraint per task, they suggest that central path approximation is a promising design principle for constrained RL algorithms. We hope this encourages further research, following this paradigm. Future directions include the extensions to high-dimensional tasks, theoretical guarantees, and applications such as safety-critical control and LLM fine-tuning.
@@ B.1 Exact Penalty Methods
We provide a general result for the exactness of the penalties considered in this work. For general discussions of exact penalty methods, we refer to standard textbooks in optimization Bertsekas [1997], Nocedal and Wright [1999]. Here, we consider a compact subset $X\subset\mathbb{R}^{n}$ with non-empty interior, differentiable functions $f,g\in C^{1}(X)$, and the constrained optimization problem
@@ B.2 C3PO Exact Penalty
and for $\mathbb{A}_{\textup{B}}\geq\mathbb{A}_{c}^{\pi_{k}}(\pi)>0$ it holds that $\delta_{\textup{B}}\geq D_{\textup{B}}(\pi,\pi_{k})>0$. To solve for $\mathbb{A}_{\textup{B}}$, we rewrite 31 as
@@
$P_{\textup{Barrier}}$ $=\{\pi:\mathbb{A}_{c}^{\pi_{k}}(\pi)-\mathbb{A}_{\textup{B}}<0,\ \mathbb{A}_{c}^{\pi_{k}}(\pi)\geq 0\},$ (35)
$=\{\pi:\mathbb{A}_{c}^{\pi_{k}}(\pi)-b\ (W(-\exp(-\delta_{\textup{B}}-1))+1)<0,\ \mathbb{A}_{c}^{\pi_{k}}(\pi)\geq 0\},$ (36)
$=\{\pi:\mathbb{A}_{c}^{\pi_{k}}(\pi)-b\ w<0,\ \mathbb{A}_{c}^{\pi_{k}}(\pi)\geq 0\},$ (37)
@@
$\max_{\pi\in\Pi}\ \mathbb{A}_{r}^{\pi_{k}}(\pi)\$$\textrm{ s.t. }\ \mathbb{A}_{c}^{\pi_{k}}(\pi)-\mathbb{A}_{\textup{B}}<0\$$\textrm{ and }\ \bar{D}_{\textup{KL}}(\pi,\pi_{k})<\delta_{\textup{KL}}$ (38)
$\max_{\pi\in\Pi}\ \mathbb{A}_{r}^{\pi_{k}}(\pi)\$$\textrm{ s.t. }\ D_{\textup{B}}(\pi,\pi_{k})<\delta_{B}\$$\textrm{ and }\ \bar{D}_{\textup{KL}}(\pi,\pi_{k})<\delta_{\textup{KL}}$ (39)
@@
$\max_{\pi\in\Pi}\ \mathbb{A}_{r}^{\pi_{k}}(\pi)$ $-\kappa_{k}\max\{0,\mathbb{A}_{c}^{\pi_{k}}(\pi)-\min(b,w\cdot b)\}\$$\textrm{ s.t. }\ \bar{D}_{\textup{KL}}(\pi,\pi_{k})<\delta_{\textup{KL}},$ (40)
$\max_{\pi\in\Pi}\ \mathbb{A}_{r}^{\pi_{k}}(\pi)$ $-\kappa_{k}\max\{0,D_{\textup{B}}(\pi,\pi_{k})-\delta_{B}\}\$$\textrm{ s.t. }\ \bar{D}_{\textup{KL}}(\pi,\pi_{k})<\delta_{\textup{KL}},$ (41)